# Compact stellar systems hosting an intermediate mass black hole: magnetohydrodynamic study of inflow-outflow dynamics

### v1

@@ 6 Conclusions
[…]
From the simulations, we found that only $10^{-5}$ of the injected wind mass reaches the black hole. Wind-wind interactions dominate the dynamics, and the system is outflow-dominated. The gas follows a virial temperature profile, reaching $T\sim 10^{8}$ K near the boundary, while high wind velocities exceed the escape speed, ensuring most gas remains unbound. In higher-metallicity environments, enhanced radiative cooling leads to the formation of cool, dense clumps, but their transient nature and high velocities prevent efficient accretion. The inflowing gas remains well below the Keplerian angular momentum, favoring a quasi-Bondi-like, radial accretion mode. Magnetic fields slightly suppress outflows but do not significantly affect the net accretion rate.
Observationally, the system’s X-ray emission is dominated by wind-wind shocks, with total luminosities of $L_{X}\sim 10^{34}-10^{35}\,\mathrm{erg\,s}^{-1}$, while the extrapolated near-horizon flow around the IMBH is much fainter at $L_{X}\sim 10^{32}\,\mathrm{erg\,s}^{-1}$. X-ray emission is wind-dominated, so IMBH signatures remain below current detection thresholds. Although weak, the X-ray emission of the near-horizon flow around the IMBH is highly variable with quasiperiodic flares recurring every $\sim 100-200$ years due to the close passages of the nearest star with the surrounding stars.
These results suggest that IMBHs in dense stellar clusters grow slowly if they are fed solely by stellar winds. In this quiescent regime, they do not affect their surroundings significantly through feedback. Instead, wind-wind collisions due to the cluster stars influence the closest environment through kinetic feedback. Feedback from the IMBH could be enhanced by, e.g. the tidal disruption of one of the cluster stars or by the accretion of an additional reservoir of external gas. Future work should investigate these possibilities as well as study the long-term cluster evolution to better understand the detectability of IMBHs and their role in stellar dynamics.
@@ Acknowledgements.
ML acknowledges the financial support of the OPUS-LAP/GAČR-LA bilateral project (2021/43/I/ST9/01352/OPUS22 and GF23-04053L). ML and TP are supported by the GAČR EXPRO grant no. GX21-13491X. MZ acknowledges the financial support of the GAČR Junior Star grant no. GM24-10599M. BR and SMR are supported by the Natural Sciences & Engineering Research Council of Canada (NSERC). BR is supported by the Canadian Space Agency (23JWGO2A01), and by a grant from the Simons Foundation (MP-SCMPS-00001470). BR acknowledges a guest researcher position at the Flatiron Institute, supported by the Simons Foundation. F.P. gratefully acknowledges the Collaborative Research Center 1601 funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – SFB 1601 [sub-project A3] – 500700252. The computational resources and services used in this work were partially provided by facilities supported by the VSC (Flemish Supercomputer Center), funded by the Research Foundation Flanders (FWO) and the Flemish Government – department EWI, and by Compute Ontario and the Digital Research Alliance of Canada (alliancecan.ca) compute allocation rrg-ripperda.
@@ Appendix A Mutual distances of stars
In this Appendix we calculate the temporal evolution of the separations of stars in our simulations, adopting Keplerian orbits based on the randomly generated orbital elements that yield the projected orbital radii consistent with observations of early-type stars in the Galactic center IRS 13 E association. We consider two orbital distributions of 6 early-type stars (see Table 3): isotropic and disk-like. In Table 4, we list the basic statistics of the separations among all the pairs of stars for the isotropic orbital distribution, including the minimum, mean, and maximum distances (expressed in astronomical units) as well as the characteristic recurrence timescale (expressed in years) based on the most prominent peaks in the Lomb-Scargle periodogram of the mutual separation. In Table 5, we list the same information for the disk-like distribution.
@@ Appendix B Periodicities and power spectral densities of the inflow rate
Figure 13: Periodicities and the power spectral density for the accretion rate in the isotropic Run I. Left: The Lomb-Scargle periodogram (power vs. period in years) with the most prominent peak at $\sim 106$ years. Other prominent peaks are at $\sim 191$ and $\sim 68$ years. Right: The PSD with the best-fit power-law function with the slope of $-0.65\pm 0.17$.
Figure 14: Periodicities and the power spectral density for the accretion rate in the isotropic Run II. Left: The Lomb-Scargle periodogram (power vs. period in years) with the most prominent peak at $\sim 112$ years. Other prominent peaks are at $\sim 204$ and $\sim 54$ years. Right: The PSD with the best-fit power-law function with the slope of $-0.57\pm 0.12$.
Figure 15: Periodicities and the power spectral density for the accretion rate in isotropic Run III. Left: The Lomb-Scargle periodogram (power vs. period in years) with the most prominent peak at $\sim 108$ years. Other prominent peaks are at $\sim 190$, $\sim 69$, and $\sim 55$ years. Right: The PSD with the best-fit power-law function with the slope of $-0.49\pm 0.20$.
Figure 16: Periodicities and the power spectral density for the accretion rate in the disk-like Run. Left: The Lomb-Scargle periodogram (power vs. period in years) with the most prominent peak at $\sim 186$ years. Other prominent peaks are at $\sim 217$, and $\sim 100$ years. Right: The PSD with the best-fit power-law function with the slope of $-0.68\pm 0.20$.

### v2

@@ 6 Conclusions
We conducted 3D MHD simulations to study wind-fed accretion onto an IMBH of mass $M_{\bullet}=3\times 10^{4}\,M_{\odot}$ embedded in a compact stellar cluster of six WR stars. Using the grid-based code athena++, we modeled the stellar winds (with wind velocities of $v_{w}\sim\,1000\ \mathrm{km\,s^{-1}}$) and their interactions with the IMBH’s gravitational potential while varying the cluster’s stellar distribution, wind magnetization, and chemical composition.
From the simulations, we found that only $10^{-5}$ of the injected wind mass reaches the black hole. Wind-wind interactions dominate the dynamics, and the system is outflow-dominated. The gas follows a virial temperature profile, reaching $T\sim 10^{8}$ K near the boundary, while high wind velocities exceed the escape speed, ensuring most gas remains unbound. In higher-metallicity environments, enhanced radiative cooling forms cool, dense clumps, but their transient nature and high velocities prevent efficient accretion. The inflowing gas remains well below the Keplerian angular momentum, favoring a quasi-Bondi-like, radial accretion mode. Magnetic fields slightly suppress outflows but do not significantly affect the net accretion rate.
Observationally, the system’s X-ray emission dominated wind-wind shocks, with total luminosities of $L_{X}\sim 10^{34}-10^{35}\,\mathrm{erg\,s}^{-1}$, while the extrapolated near-horizon flow around the IMBH is much fainter at $L_{X}\sim 10^{32}\,\mathrm{erg\,s}^{-1}$. X-ray emission is wind-dominated, so IMBH signatures remain below current detection thresholds. Although weak, the X-ray emission of the near-horizon flow around the IMBH is highly variable with quasiperiodic flares recurring every $\sim 100-200$ years due to the close passages of the nearest star with the surrounding stars.
These results suggest that IMBHs in dense stellar clusters grow slowly if fed solely by stellar winds. In this quiescent regime, they do not affect their surroundings significantly through feedback. Instead, wind-wind collisions due to the cluster stars influence the closest environment through kinetic feedback. Feedback from the IMBH could be enhanced by, e.g., the tidal disruption of one of the cluster stars or the accretion of an additional external gas reservoir. Future work should investigate these possibilities and study the long-term cluster evolution to better understand the detectability of IMBHs and their role in stellar dynamics.
@@ Acknowledgements.
We thank S. von Fellenberg for useful discussions. ML and TP have received support from the GAČR EXPRO grant no. GX21-13491X. MZ acknowledges the financial support of the GAČR Junior Star grant no. GM24-10599M. BR and SMR supported the Natural Sciences & Engineering Research Council of Canada (NSERC). BR supported the Canadian Space Agency (23JWGO2A01), and by a grant from the Simons Foundation (MP-SCMPS-00001470). BR acknowledges a guest researcher position at the Flatiron Institute, supported by the Simons Foundation. F.P. gratefully acknowledges the Collaborative Research Center 1601 funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – SFB 1601 [sub-project A3] – 500700252. The computational resources and services used in this work were partially provided by facilities supported by the VSC (Flemish Supercomputer Center), funded by the Research Foundation Flanders (FWO) and the Flemish Government – department EWI, and by Compute Ontario and the Digital Research Alliance of Canada (alliancecan.ca) compute allocation rrg-ripperda.
@@ Appendix A Mutual distances of stars
In this Appendix, we calculate the temporal evolution of the separations of stars in our simulations, adopting Keplerian orbits based on the randomly generated orbital elements that yield the projected orbital radii consistent with observations of early-type stars in the Galactic center IRS 13 E association. We consider two orbital distributions of 6 early-type stars (see Table 3): isotropic and disk-like. In Table 4, we list the basic statistics of the separations among all the pairs of stars for the isotropic orbital distribution, including the minimum, mean, and maximum distances (expressed in astronomical units) as well as the characteristic recurrence timescale (expressed in years) based on the most prominent peaks in the Lomb-Scargle periodogram of the mutual separation. In Table 5, we list the same information for the disk-like distribution.
@@ Appendix B Periodicities and power spectral densities of the inflow rate
Figure 13: Periodicities and the power spectral density for the accretion rate in the isotropic Run I. Left: The Lomb-Scargle periodogram (power vs. period in years) with the most prominent peak at $\sim 106$ years. Other prominent peaks are at $\sim 191$ and $\sim 68$ years. Right: The PSD with the best-fit power-law function with the slope of $0.65\pm 0.17$.
Figure 14: Periodicities and the power spectral density for the accretion rate in the isotropic Run II. Left: The Lomb-Scargle periodogram (power vs. period in years) with the most prominent peak at $\sim 112$ years. Other prominent peaks are at $\sim 204$ and $\sim 54$ years. Right: The PSD with the best-fit power-law function with the slope of $0.57\pm 0.12$.
Figure 15: Periodicities and the power spectral density for the accretion rate in isotropic Run III. Left: The Lomb-Scargle periodogram (power vs. period in years) with the most prominent peak at $\sim 108$ years. Other prominent peaks are at $\sim 190$, $\sim 69$, and $\sim 55$ years. Right: The PSD with the best-fit power-law function with the slope of $0.49\pm 0.20$.
Figure 16: Periodicities and the power spectral density for the accretion rate in the disk-like Run. Left: The Lomb-Scargle periodogram (power vs. period in years) with the most prominent peak at $\sim 186$ years. Other prominent peaks are at $\sim 217$, and $\sim 100$ years. Right: The PSD with the best-fit power-law function with the slope of $0.68\pm 0.20$.
Appendix C Comparison of the IRS 13E X-ray surface-brightness profile with the mock cluster
In this Appendix, we compare the IRS 13E X-ray surface-brightness profile with the one inferred from simulations. In particular, we consider the profiles in two directions: East-West ($x$) and North-South ($y$). In addition, we extract the profile of the Chandra point-spread function (PSF). In Fig. 17, we compare the xyx-y surface-brightness profiles for the observed Chandra image as well as for the simulated one with the Chandra PSF. We see that the observed profile is comparable to the PSF close to the peak, while it is slightly extended towards the tails. The profile constructed from the simulated X-ray image of the compact cluster is consistent with the observed one, except for the wing to the west, where the observed profile is significantly more extended. This can be attributed to the surrounding diffuse X-ray emission and/or individual X-ray sources. This particular “bump” in the western tail (positive $x$-direction) was previously attributed to the source E60 (Wang et al. 2020; Zhu et al. 2020).
In terms of the image asymmetry, we compare the observed IRS 13E image with the simulated one using the asymmetry index defined as follows,
$A=\frac{\sum_{i}|I_{i}-I_{{\rm rot},i}|}{2\sum_{i}|I_{i}|}$ (12)
where we subtract the image rotated by $180^{\circ}$ from the actual image. The observed image has $A=0.21$ while the simulated one has $A=0.13$, which is consistent with the simulated image being more symmetric as is also visible in Fig. 17. As we already mentioned, in the observed image, there is a contribution at least from the source E60, which is not a part of the IRS 13E association and as such is not included in the simulated cluster.
Figure 17: Comparison of the observed IRS 13E X-ray surface-brightness profiles (black and red points with errorbars in normalized units in two perpendicular directions $x$ and $y$ expressed in pixels) with the corresponding ones inferred from simulations presented in this study (black and red dashed lines). A dotted blue line depicts the Chandra PSF. The more extended western wing of IRS 13E can be attributed to the E60 source. One pixel is equivalent to $\sim 0.125$ arcsec or $\sim 5\times 10^{-3}\,{\rm pc}$ at the Galactic center.